OpenStudy (ineedhelpfast12):

solve the linear system using substitution. 2x=12 x-5y=-29

OpenStudy (susanna):

first divide both sides by 2 in the equation 2x=12

OpenStudy (susanna):

x=6

OpenStudy (susanna):

now, substitute 6 for x in the second equation. (6)-5y=-29

OpenStudy (susanna):

Subtract 6 from both sides: -5y=-35

OpenStudy (susanna):

Divide both sides by -5 y=7

OpenStudy (susanna):

Plug in 7 for y to get the x x-5(7)=-29 x-35=-29 x=6

OpenStudy (ineedhelpfast12):

y=7 x=6

OpenStudy (susanna):

That was a check, because from the first equation you already know that x=6.

OpenStudy (ineedhelpfast12):

thank you very much

OpenStudy (ineedhelpfast12):

can you help me with one more?

OpenStudy (susanna):

sure

OpenStudy (susanna):

make a new post